Skid mark analysis is one way to estimate the speed a car was traveling prior to an accident. The speed, s, that the vehicle was traveling in miles per hour can be approximated by the formula s=square root 30df, where d represents the length of the skid marks in fee and "f" represents the drag factor of the road. A vehicle that was involved in an accident on an asphalt road was traveling at a speed of 45 mph when it started skidding. If the drag factor for asphalt is 0.75 find the length of the skid marks made by the car. The skid marks made by the car were approximately___________ feet long.
Answer by stanbon(75887)

s=square root 30df, where d represents the length of the skid marks in feet and "f" represents the drag factor of the road. A vehicle that was involved in an accident on an asphalt road was traveling at a speed of 45 mph when it started skidding. If the drag factor for asphalt is 0.75 find the length of the skid marks made by the car.
S = sqrt(30df)
45 = sqrt(30*d*0.75)
----
45 = sqrt[(90/4)d]
-----
45 = (3/2)sqrt(10d)
30 = sqrt(10d)
10d = 900
d = 90 ft.
================
Cheers,
Stan H.





The skid marks made by the car were approximately___________ feet long.